OpenStudy (anonymous):

logb(x+2)−logb4=logb3x solve for x

OpenStudy (anonymous):

log (x+2)/3x =log 4 (x+2)/3x =4

OpenStudy (anonymous):

now some algebra

OpenStudy (anonymous):

x=2/11

OpenStudy (anonymous):

x= -2/11